6. It requires 222 kJ/mol to eject electrons from potassium metal. photoelectric effect? (b) What is the wavelength of this light? (c) If potassium is irradiated with light of 350 nm, what is the maximum possible kinetic energy of the emitted electrons? How many orbitals are represented in an atom by the following designations? (a) p (b) 4d 7. (c) 2s (d) 5f (e) 2d
Answer #1

6.a)You need to convert from the quantity of energy need to eject 1 mole electrons to amount of energy required to eject a single electron

Therefore ,E0 = (222kJ/mole)(1000J/KJ)(1 mole/6.022 X 1023 ) =3.69 X 10-19 J

According to planks equation E=h v

v = E/ h =3.69 X 10-19 J / 6.626X 10 -34 J-S

=5.56 X 10 14 Hz or s-1

b)wavelenth of light is (λ) =C / v = (3X108 m/s) X (5.56X 10-14 s -1)

=539nm

C)If K is irradiated with 350nm,then maximum possible kinetic energy is

Ekinetic \RightarrowEpht - E0

Epht = hc / (λ) =[6.626 x 10-34 J-s][3.00 X 108 m/s] / [ 350 x 10-9 m]

=5.68 x 10-19 J

Ekinetic=  5.68X10-19 J - 3.69X 10-19 J

  \Rightarrow1.98 X10-19 J

therefore the maximum kinetic energy to emit electron is 1.98 x 10-19 J

7.a)The orbitals present in px is 1

b)The orbitals present in 4d is 5

c)The orbitals present in 2s is 1

d)The orbitals present in 5f is 7

e)The orbitals present in 2d is 5